Determine the inductance of a solenoid with 670 turns in a length of 30cm . The circular cross section of the solenoid has a radius of 5.0cm .
Answer #1

self inductance of a solenoid is L=(uo*N^2*A)/l


=(12.56*10^-7*670^2*(pi*0.05^2)/0.3

=14.75*10^-3 Henry

=14.75 mH